Luminous intensity = 64 cd

Distance between source and surface = 3.0 m

Calculate illuminace


1
Expert's answer
2021-09-18T14:59:35-0400

The illuminance is


"E=\frac{I}{R^2}=\frac{64}{9^2}=0.79\text{ lux}."